# From Estimands to Robust Inference of Treatment Effects in Platform Trials

### v1

@@ 5.2 Efficiency comparison
Figure 3: Efficiency comparisons among all robust estimators.
6 Robust Variance Estimation
To robustly assess variability and conduct large sample inference, it is essential to obtain consistent estimators of the asymptotic covariance matrices in Theorem 1 for various estimators of $\bm{\vartheta}_{jk}$. These estimators should remain consistent whether the working models are correct or misspecified.
For IPW-type estimators, we propose the following estimators of covariance matrices:
$\hat{\bm{\Sigma}}_{jk}^{\rm\,ipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{\cal I}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)Y_{i}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)Y_{i}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}-\hat{\bm{\vartheta}}_{jk}\hat{\bm{\vartheta}}_{jk}^{T},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,sipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)(Y_{i}-\hat{\theta}_{jk})^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)(Y_{i}-\hat{\theta}_{kj})^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk}-\hat{\bm{\delta}}_{jk}\hat{\bm{\delta}}_{jk}^{T},$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,saipw}\!=\!\frac{1}{n_{jk}}\sum_{i\in{{{\cal I}}}_{jk}}\!{\rm diag}\!\left\{{\textstyle\frac{I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})-\hat{\delta}_{jk}\}^{2}}{\pi_{j}^{2}({Z}_{i})},\,\frac{I(A_{i}=k)\{Y_{i}-\hat{\mu}_{kj}(\bm{X}_{i})-\hat{\delta}_{kj}\}^{2}}{\pi_{k}^{2}({Z}_{i})}}\right\}+\hat{\bm{\Lambda}}_{jk},$
where $\hat{\bm{\vartheta}}_{jk}=(\hat{\theta}_{jk},\hat{\theta}_{kj})^{T}$, $\hat{\theta}_{jk}$ is any estimator of $\theta_{jk}$, $\hat{\bm{\delta}}_{jk}=(\hat{\delta}_{jk},\hat{\delta}_{kj})^{T}$, $\hat{\delta}_{jk}=n_{jk}^{-1}\sum_{i\in{{{\cal I}}}_{jk}}\!I(A_{i}=j)\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}/\pi_{j}({Z}_{i})$, $\hat{\bm{\Lambda}}_{jk}$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ respectively estimated by $\hat{\lambda}_{jk}=2\hat{q}_{jk}^{(j)}+\hat{\sigma}^{2}_{jk}$ and $\hat{c}_{jk}=\hat{q}_{kj}^{(j)}+\hat{q}_{jk}^{(k)}+\hat{q}_{jk}$, $\hat{q}_{jk}^{(j)}$ is the sample covariance of $(Y_{i}^{(j)}-\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{jk}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$ and $A_{i}=j$, $\hat{q}_{kj}^{(j)}$ is $\hat{q}_{jk}^{(j)}$ with $\hat{\mu}_{jk}(\bm{X}_{i})$ switched to $\hat{\mu}_{kj}(\bm{X}_{i})$, $\hat{q}_{jk}^{(k)}$ is $\hat{q}_{jk}^{(j)}$ with $Y_{i}^{(j)}$ switched to $Y_{i}^{(k)}$, $\hat{\sigma}^{2}_{jk}$ and $\hat{\sigma}^{2}_{kj}$ are the diagonal entries and $\hat{q}_{jk}$ is the off-diagonal entry of the sample covariance matrix of $(\hat{\mu}_{jk}(\bm{X}_{i}),\hat{\mu}_{kj}(\bm{X}_{i}))$’s with $i\in{\cal I}_{jk}$.
For the PS and APS estimators, we propose the following estimators of $\bm{\Sigma}_{jk}^{\rm\,ps}$ and $\bm{\Sigma}_{jk}^{\rm\,aps}$,
$\hat{\bm{\Sigma}}_{jk}^{\rm\,ps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,{\rm diag}\!\left\{{\textstyle\frac{\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\sigma}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Gamma}}_{jk}$
$\hat{\bm{\Sigma}}_{jk}^{\rm\,aps}$ $=\sum_{h=1}^{H_{jk}}\frac{n({\cal I}_{jk}^{(h)})}{n_{jk}}\,\Big{[}{\rm diag}\!\left\{{\textstyle\frac{\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{j}({\cal I}_{jk}^{(h)})},\,\frac{\hat{\tau}_{k}^{2}({\cal I}_{jk}^{(h)})}{\hat{\pi}_{k}({\cal I}_{jk}^{(h)})}}\right\}+\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})\Big{]}+\hat{\bm{\Gamma}}_{jk},$
where $\hat{\pi}_{j}({\cal I}_{jk}^{(h)})=n_{j}({\cal I}_{jk}^{(h)})/n({\cal I}_{jk}^{(h)})$, $\hat{\sigma}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $Y_{i}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\tau}_{j}^{2}({\cal I}_{jk}^{(h)})$ is the sample variance of $\{Y_{i}-\hat{\mu}_{jk}(\bm{X}_{i})\}$’s with $i\in{\cal I}_{jk}^{(h)}$ and $A_{i}=j$, $\hat{\bm{\Lambda}}_{jk}({\cal I}_{jk}^{(h)})$ is $\bm{\Lambda}_{jk}$ with $\lambda_{jk}$ and $c_{jk}$ estimated respectively by $\hat{\lambda}_{jk}({\cal I}_{jk}^{(h)})=2\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})+\hat{\sigma}_{jk}^{2}({\cal I}_{jk}^{(h)})$ and $\hat{c}_{jk}({\cal I}_{jk}^{(h)})=\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}^{(k)}({\cal I}_{jk}^{(h)})+\hat{q}_{jk}({\cal I}_{jk}^{(h)})$, $\hat{q}_{jk}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{q}_{kj}^{(j)}({\cal I}_{jk}^{(h)})$, $\hat{\sigma}^{2}_{jk}({\cal I}_{jk}^{(h)})$, and $\hat{q}_{jk}({\cal I}_{jk}^{(h)})$ are respectively the same as $\hat{q}_{jk}^{(j)}$, $\hat{q}_{kj}^{(j)}$, $\hat{\sigma}^{2}_{jk}$, and $\hat{q}_{jk}$ previously defined in $\hat{\bm{\Sigma}}_{jk}^{\rm\,aipw}$ but with the index $i$ restricted to $i\in{\cal I}_{jk}^{(h)}$, $\hat{\bm{\Gamma}}_{jk}$ is the sample covariance matrix of $(\bar{Y}_{j}(S_{i}),\bar{Y}_{k}(S_{i}))$’s for $i\in{\cal I}_{jk}$, and $\bar{Y}_{j}(S_{i})$ is the sample mean of $Y_{i}$’s with $i\in{\cal I}_{jk}$, $A_{i}=j$, and $S_{i}={\cal I}_{jk}^{(h)}$.
7 Empirical Results
7.1 Simulation study
We conduct a simulation study based on the first three enrollment windows outlined in Figure 1 to examine the finite-sample performance of all estimators presented in Section 4.
In the simulation, the observed baseline covariate vector is $\bm{W}=(X_{c},X_{b},Z_{\rm sub},Z_{\rm win})$, where $X_{c}$ is a continuous covariate uniformly distributed over the interval $(-3,3)$, $X_{b}$ is a binary covariate with $P(X_{b}=1)=0.5$, $Z_{\rm sub}$ is a binary indicator for disease subtype with $P(Z_{\rm sub}=1)=0.8$, and $Z_{\rm win}\in\{1,2,3\}$ denotes the enrollment window and satisfies
$P(Z_{\rm win}=t\mid X_{c},X_{b},Z_{\rm sub},U)=\frac{\exp(Q_{t})}{\exp(Q_{1})+\exp(Q_{2})+\exp(Q_{3})},\quad t=1,2,3,$
$Q_{1}=0.5+X_{c}+2X_{b}-Z_{\rm sub}+U$, $Q_{2}=1+2X_{c}+X_{b}-Z_{\rm sub}+U$, $Q_{3}=-0.5+X_{c}+X_{b}+Z_{\rm sub}+U$, and the unobserved baseline variable $U$ follows a standard normal distribution and is independent of $X_{c}$, $X_{b}$, and $Z_{\rm sub}$. The potential outcomes are generated as follows: $Y^{(1)}=1+X_{c}+X_{b}+Z_{\rm sub}+U+\epsilon_{1},$ $Y^{(2)}=1+X_{c}^{2}+X_{b}+Z_{\rm sub}+U+\epsilon_{2}$, $Y^{(3)}=3+X_{c}X_{b}+Z_{\rm sub}+U+\epsilon_{3}$, $Y^{(4)}=2+X_{c}Z_{\rm sub}-X_{b}+2U+\epsilon_{4}$, where $\epsilon_{j}$’s are mutually independent, each follows a standard normal distribution and is independent of $\bm{W}$ and $U$. Note that $Y^{(j)}$’s are correlated with $X_{c}$, $X_{b}$, $Z_{\rm sub}$, $U$, as well as indirectly with the enrollment time $Z_{\rm win}$.
Table 1: Assignment probabilities to each sub-study and treatment
${Z}$ Sub-study Treatment
$Z_{\rm win}$ $Z_{\rm sub}$ 1 2 3 $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$ $\pi_{4}({Z})$
1 1 0.4 0.6 0 0.50 0.20 0.30 0
1 0 1 0 0 0.50 0.50 0 0
2 1 0.3 0.3 0.4 0.50 0.15 0.15 0.20
2 0 1 0 0 0.50 0.50 0 0
3 1 0.4 0 0.6 0.50 0.20 0 0.30
3 0 1 0 0 0.50 0.50 0 0
Patients are first randomized into one of three sub-studies (labeled 1-3) depending on their enrollment windows and disease subtypes, and then to treatments within the assigned sub-study. Sub-study 1 has treatments 1 and 2, sub-study 2 has treatments 1 and 3, and sub-study 3 has treatments 1 and 4. Within each sub-study, treatments are randomized in a 1:1 ratio. The assignment probabilities are in Table 1. For the total sample size $n$ in the platform trial, we consider 500 and 1,000. When $n=500$, the expected numbers of patients per treatment are 123, 51, and 76 in sub-studies 1-3, respectively.
We consider the following estimators of the linear contrast $\theta_{j1}-\theta_{1j}$, for $j=2,3,4.$
  1. (i)
    Four IPW-type estimators. We consider the IPW, SIPW, and SAIPW estimators described in Section 4.2. For SAIPW, $\hat{\mu}_{jk}$ is obtained from fitting a linear regression of $Y$ on $(X_{c},X_{b},Z_{\rm sub})$, using data from treatment arm $j$ in ${\cal I}_{jk}$. Since condition (4) holds in this simulation, the AIPW estimator is numerically almost the same as the SAIPW estimator and thus is omitted. We add the SAIPW estimator with $\bm{X}=S$, denoted by SAIPW($S$), to check its equivalence with the PS estimator as discussed in Corollary 5. Note that the working models are correctly specified only for treatment arm 1, as the true models for the other arms are not linear in $(X_{c},X_{b},Z_{\rm sub})$.
  2. (ii)
    Four PS-type estimators. We consider the following strata for PS and APS:
    Estimating ${\small\begin{array}[]{ll}\mbox{Estimating $\theta_{21}-\theta_{12}$:}&{\cal I}_{12}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.5\,(Z_{{\rm sub},i}=0)\},\\ &{\cal I}_{12}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.2\,(Z_{{\rm win},i}=1\ \mbox{or}\ 3,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{12}^{(3)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{2}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{31}-\theta_{13}$:}&{\cal I}_{13}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.3\,(Z_{{\rm win},i}=1,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{13}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{3}({Z}_{i})=0.15\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\};\\ \mbox{Estimating $\theta_{41}-\theta_{14}$:}&{\cal I}_{14}^{(1)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.2\,(Z_{{\rm win},i}=2,Z_{{\rm sub},i}=1)\},\\ &{\cal I}_{14}^{(2)}=\{i:\pi_{1}({Z}_{i})=0.5,\pi_{4}({Z}_{i})=0.3\,(Z_{{\rm win},i}=3,Z_{{\rm sub},i}=1)\}.\end{array}}\vspace{2mm}$
    The same working models $\hat{\mu}_{jk}$ for SAIPW are used for APS. For estimating $\theta_{21}-\theta_{12}$, we additionally include the PS and APS estimators that stratify on all six joint levels of ${Z}=(Z_{\rm win},Z_{\rm sub})$, denoted as PS($Z$) and APS($Z$). For estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$, we have PS($Z$) = PS and APS($Z$) = APS because using all six joint levels of ${Z}$ results in the same two strata as above (due to $\pi_{j}(Z)=0$ for some $j=3,4$).
  3. (iii)
    The naive estimator. For comparison we include the naive estimator of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, as described in Section 4.1.
  4. (iv)
    Estimators within each sub-study. For comparison, we also evaluate traditional ANOVA (sample means) and ANCOVA estimators (using linear working models to adjust for $X_{c},X_{b}$, and $Z_{\rm sub}$) based solely on data from each individual sub-study. Although these sub-study-specific estimators estimate the treatment effect within each sub-study, which differs from our target estimand defined in (1), we can still compare their variances.
The simulation results based on $5,000$ runs are reported in Table 2 for the previously described eleven estimators in (i)-(iv). The results contain the average bias, standard deviation (SD), the average of the standard error (SE), and the coverage probability (CP) of the $95\%$ confidence interval. The true values of estimands are approximated using an independent simulated dataset with size $10^{7}$. The following is a summary of the simulation results in Table 2.
Table 2: Simulation results based on $5,000$ simulation runs
$\theta_{21}-\theta_{12}=3$ $\theta_{31}-\theta_{13}=1.145$ $\theta_{41}-\theta_{14}=-0.886$
$n$ Method Bias SD SE CP Bias SD SE CP Bias SD SE CP
500 Naive -0.231 0.320 0.316 0.874 -0.185 0.342 0.340 0.916 -0.205 0.384 0.380 0.911
IPW -0.006 0.639 0.636 0.946 0.004 0.776 0.777 0.948 -0.007 0.500 0.497 0.948
SIPW -0.003 0.341 0.336 0.941 0.005 0.347 0.341 0.943 0.001 0.389 0.381 0.942
SAIPW -0.018 0.329 0.340 0.951 0.001 0.284 0.284 0.944 -0.001 0.297 0.300 0.949
$(S)$ 0.000 0.336 0.329 0.939 0.009 0.327 0.324 0.939 0.002 0.356 0.348 0.942
PS 0.000 0.336 0.335 0.945 0.009 0.327 0.330 0.949 0.002 0.356 0.356 0.946
APS -0.013 0.329 0.339 0.952 -0.001 0.286 0.289 0.947 -0.002 0.298 0.306 0.956
PS$(Z)$ Estimate or its SE cannot
APS$(Z)$ be computed in 408 runs
ANOVA* 0.001 0.354 0.350 0.946 0.006 0.421 0.417 0.945 0.004 0.424 0.425 0.946
ANCOVA* -0.008 0.329 0.321 0.942 -0.005 0.319 0.311 0.942 0.002 0.321 0.314 0.942
1000 Naive -0.230 0.226 0.224 0.819 -0.189 0.240 0.239 0.872 -0.206 0.269 0.268 0.876
IPW -0.001 0.453 0.451 0.947 0.012 0.550 0.550 0.951 0.003 0.355 0.352 0.943
SIPW 0.000 0.243 0.239 0.945 0.004 0.246 0.243 0.944 0.001 0.272 0.270 0.948
SAIPW -0.009 0.232 0.242 0.954 0.004 0.198 0.203 0.955 0.000 0.212 0.213 0.947
$(S)$ 0.001 0.238 0.234 0.945 0.004 0.233 0.230 0.942 0.003 0.252 0.247 0.943
PS 0.001 0.238 0.236 0.948 0.004 0.233 0.232 0.944 0.003 0.252 0.250 0.947
APS -0.006 0.232 0.239 0.952 0.003 0.198 0.203 0.955 0.000 0.213 0.215 0.952
PS$(Z)$ 0.001 0.238 0.236 0.948
APS$(Z)$0.005 0.233 0.229 0.942
ANOVA* 0.003 0.251 0.248 0.947 0.007 0.295 0.294 0.944 0.002 0.301 0.299 0.948
ANCOVA* -0.003 0.233 0.228 0.947 0.005 0.222 0.220 0.947 0.001 0.227 0.222 0.945
SAIPW($S$): SAIPW with $\bm{X}=S$.
PS($Z$) and APS($Z$): post-stratification using all joint levels of $Z=(Z_{\rm win},Z_{\rm sub})$ as strata (PS = PS($Z$) and APS = APS($Z$) for estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$).
*ANOVA and ANCOVA based on data within each sub-study, estimating the within sub-study treatment effects, which are $3.054$, $1.279$, and $-0.881$ for sub-studies 1-3, respectively.
  1. 1.
    The simulation results support our asymptotic theory in Section 5. Specifically, the naive estimator is biased, while all IPW-type and PS-type estimators have negligible biases; despite the misspecification of working models, covariate adjustment improves efficiency as reflected by reductions in SD; the PS and SAIPW($S$) estimators demonstrate nearly identical performance. Additionally, SEs are close to SDs and CPs are close to the nominal level 95%.
  2. 2.
    No results are reported in Table 1 for PS($Z$) and APS($Z$) at $n=500$ because either the estimate of $\theta_{jk}$ or its SE cannot be computed in 408 out of 5,000 simulation runs. This issue arises due to insufficient data in the stratum defined by $(Z_{\rm win},Z_{\rm sub})=(3,0)$, where the average number of patients is 10. A similar issue occurs in 3 out of 5,000 runs at $n=1,000$, so the results in Table 1 are based on 4,997 runs. This problem does not affect the PS and APS estimators: our proposed stratification method combines three strata $(Z_{\rm win},Z_{\rm sub})=(1,0),(2,0),(3,0)$ into ${\cal I}_{12}^{(1)}$ because they all have constant $\pi_{1}(Z)=0.5$ and $\pi_{2}(Z)=0.5$.
  3. 3.
    Between IPW and SIPW, it is clear that SIPW is much better in terms of SD under our simulation setting. Notably, IPW can have an even larger SD than ANOVA within substudies, which usually does not occur with SIPW.
  4. 4.
    The traditional ANOVA and ANCOVA estimators analyze the platform trials as if they were separate, stand-alone studies. Without covariate adjustment, the ANOVA estimator has a much larger SD than the SIPW estimator when estimating $\theta_{31}-\theta_{13}$ and $\theta_{41}-\theta_{14}$. A similar conclusion applies to the ANCOVA estimator when compared with the SAIPW or APS estimator. The efficiency difference is not as pronounced for estimating $\theta_{21}-\theta_{12}$, although the ANOVA and ANCOVA estimators still have larger SDs than the corresponding stabilized IPW-type and PS-type estimators. The reason for this is because the traditional ANOVA and ANCOVA do not utilize data across sub-studies, for example, when $n=500$, about 123, 51, and 76 patients are used under treatment arm 1 in sub-studies 1-3, respectively, whereas about 250, 121, and 150 patients in treatment arm 1 under platform trial are shared for our proposed estimators of $\theta_{j1}-\theta_{1j}$, $j=2,3,4$, respectively.
7.2 Real data analysis
As reviewed in Section 2.2 and shown in Figure 2, the SIMPLIFY design (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study. Let $Z$ be the indicator of whether a participant receives HS or DA therapy at baseline and the enrollment window, with the assignment probabilities given in Table 3. The enrollment window (EW) is treated as a categorical variable, with EW1 representing the enrollment period before the probability change, and EW2 standing for the enrollment period after the change, as presented in Figure 2. The outcome $Y$ is the 6-week mean absolute change in ppFEV1 as described in Section 2.2.
In our analysis, the treatment effects of interest are $\theta_{21}-\theta_{12}$ and $\theta_{31}-\theta_{13}$. Specifically, $\theta_{21}-\theta_{12}$ is the effect of discontinuing HS among individuals taking HS, regardless of their DA status. This represents the treatment effect among participants with $\pi_{1}(Z)>0$ and $\pi_{2}(Z)>0$, pooled across studies (i.e., participants with $\text{HS}=1$ in Table 3). The interpretation applies similarly to $\theta_{31}-\theta_{13}$. Note that our estimand is slightly different from the estimand in Mayer-Hamblett et al., (2023), as the latter analyzed the HS and DA studies separately and included participants who re-enrolled for an additional study as noted in Section 2.2. We also excluded 10 (1.7%) participants due to missing outcomes, resulting in a sample size of $n=584$: 293 in treatment 1, 119 in treatment 2, and 172 in treatment 3.
Table 3: Assignment probabilities in the SIMPLIFY trial
${Z}$ Study Treatment
HS DA EW HS DA $\pi_{1}({Z})$ $\pi_{2}({Z})$ $\pi_{3}({Z})$
1 1 EW1 0.50 0.50 0.50 0.25 0.25
1 1 EW2 0.75 0.25 0.50 0.375 0.125
1 0 any 1 0 0.50 0.50 0
0 1 any 0 1 0.50 0 0.50
We consider the IPW, SIPW, SAIPW, PS, and APS estimators, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section 7.1. For SAIPW and APS, in addition to $Z$, we adjust several other baseline covariates, including two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, delta F508 heterozygous, and other or unknown).
Estimates and 95% confidence intervals are presented in Figure 4, where all values are given in %. Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 4), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA. The conclusion is consistent with those in Mayer-Hamblett et al., (2023).
Figure 4: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial

### v4

@@ 5.2 Efficiency comparison
Figure 5: Efficiency comparisons among all robust estimators, with the corresponding corollary annotated for each comparison.
Lastly, we discuss the data used by the estimation methods as an informal way to conceptualize the efficiency comparison. When comparing treatments $j$ and $k$, all methods can robustly utilize the outcome data from arms $j$ and $k$ within ${{{\cal I}}}_{jk}$, which may achieve substantial efficiency gains compared to traditional stand-alone trials. Furthermore, covariate adjustment methods (including PS, APS, AIPW, and SAIPW) can leverage additional strata and covariate information from all individuals within ${{{\cal I}}}_{jk}$, even those assigned to arms other than $j$ and $k$. This can lead to further efficiency gains, especially in platform trials with many arms.
6 Results
As reviewed in Section 2.1 and shown in Figure 1, the SIMPLIFY trial (Mayer-Hamblett et al.,, 2023) consists of two sub-studies (the HS and DA studies) and a total of three treatments: continuing therapy (treatment 1, used as the control) in both the HS and DA studies, discontinuing HS (treatment 2) in the HS study, and discontinuing DA (treatment 3) in the DA study. After completing the first study, participants on both HS and DA therapies may re-enroll in the other study contingent on eligibility and willingness to consent. The outcome $Y$ is the 6-week mean absolute change in ppFEV1 as described in Section 2.2.
In our analysis, the ECE population for evaluating the effect of discontinuing HS versus continuing is the population of all individuals taking only HS or both HS and DA. Accordingly, the treatment effect of interest, $\theta_{21}-\theta_{12}$, is defined on this ECE population. A similar interpretation applies to $\theta_{31}-\theta_{13}$. Demographic and baseline characteristics for participants in the two ECE populations are presented in Tables S1-S2 of the Supplement. Note that our estimand differ slightly from those in Mayer-Hamblett et al., (2023). They analyzed the HS and DA studies separately according to their primary hypotheses of interest and included data collected after participants re-enrolled in an additional study. In contrast, we excluded data recorded after re-enrollment, as our methods currently do not account for repeated measures. We plan to address this in future research. We also excluded 10 (1.7%) participants due to missing outcomes, resulting in a sample size of $n=584$: 293 in treatment 1, 119 in treatment 2, and 172 in treatment 3.
For the defined estimands, we consider the IPW, SIPW, SAIPW, PS, and APS estimators, as described in Section 4, all of which can account for the different treatment assignment probabilities across $Z$ and are evaluated in the simulation study in Section S2 of the Supplement. Here, $Z$ includes baseline use of HS, DA, and enrollment windows. For SAIPW and APS, $\hat{\mu}_{jk}$ is estimated by fitting a linear regression of $Y$ on covariates using data from treatment arm $j$ in $\mathcal{I}_{jk}$. The covariates include $Z$ and several other baseline variables: two continuous covariates, age and baseline ppFEV1; three baseline binary covariates, sex, race (white or non-white), and ethnicity (Hispanic or Latino, or not Hispanic or Latino); and one three-category covariate: genotype (delta F508 homozygous, heterozygous, or other/unknown). As shown in Corollary 2 and simulations, when adjusting for covariates using linear working models, the AIPW estimator is numerically nearly identical to the SAIPW estimator and thus omitted. We also include the naive ANOVA as described in Section 4.2, which is subject to confounding bias. Results are shown in Figure 6.
Figure 6: Estimates and 95% confidence intervals (in %) in the SIMPLIFY trial. Naive refers to the naive ANOVA method described in Section 4.2. ANOVA* and ANCOVA* are estimates based on data from each individual sub-study. IPW, SIPW, SAIPW, PS, and APS are the correct estimators for the estimand defined on the ECE population.
For all five proposed estimators (IPW, SIPW, SAIPW, PS, and APS), all point estimates are similar, as they target the same estimand defined on the ECE population. Since all confidence intervals are on the right side of the pre-defined non-inferiority margin --3% (the vertical dashed line in Figure 6), non-inferiority of discontinuing therapy can be clearly claimed for both HS and DA – consistent with the findings reported in Mayer-Hamblett et al., (2023). IPW and SIPW produce similar confidence intervals, likely because the true effect is near zero in this non-inferiority study. SIPW and SAIPW, as well as PS and APS, also yield similar confidence intervals, as most covariates are only weakly associated with the outcome ($R^{2}$ between 0.023 and 0.1 across three arms). This likely also explains why the naive estimator does not exhibit substantial bias in this analysis.
For comparison, we also include the ANOVA* (sample means) and ANCOVA* estimators (using linear working models to adjust for the same covariates included in SAIPW and APS separately for each arm) based solely on data from each individual sub-study. As discussed in Section 3.2, these estimators target analysis populations assigned to the HS or DA study. As a result, the interpretations of these estimators differ from the five proposed estimators. Specifically, for the HS comparison, the HS study includes 18 HS-only and 222 HS+DA participants, while the sample from the ECE population includes 18 HS-only and 427 HS+DA participants. For the DA comparison, the DA study includes 139 DA-only and 205 HS+DA participants, whereas the sample from the ECE population includes 139 DA-only and 427 HS+DA participants. More detailed demographic and baseline characteristics for participants in the HS and DA studies are presented in Tables S3 and S4. The larger population difference in the DA comparison likely explains the more noticeable discrepancy in point estimates for that result (opposite sign relative to the five proposed estimators).
Moreover, as expected, the ANOVA* and ANCOVA* estimators yield wider confidence intervals compared to the five proposed estimators. This is because, for the five proposed estimators, participants on both HS and DA assigned to treatment 1 ($n=214$) contribute to both the treatment 2 vs. 1 and treatment 3 vs. 1 comparisons. The covariate-adjusted methods (SAIPW and APS) further leverage covariate information from all participants on both HS and DA ($n=427$) in both comparisons. In contrast, ANOVA* and ANCOVA* split these participants between the HS and DA sub-studies, reducing the sample size for each comparison. As a result, the five proposed estimators achieve approximately 5%–27% variance reduction relative to the sub-study-specific estimators.
Overall, the results from the SIMPLIFY trial highlight the critical importance of clearly defining the estimand in umbrella and platform trials. Without this clarity, estimators may have interpretations that depend on the randomization ratio or trial design. An additional advantage of the five proposed estimators is the potential for efficiency gains through robust use of pooled data across sub-studies.